### v2

# Synergy over Discrepancy: A Partition-Based Approach to Multi-Domain LLM Fine-Tuning

Abstract
Large language models (LLMs) demonstrate impressive generalization abilities, yet adapting them effectively across multiple heterogeneous domains remains challenging due to inter-domain interference. To overcome this challenge, we propose a partition-based multi-stage fine-tuning framework designed to exploit inter-domain synergies while minimizing negative transfer. Our approach strategically partitions domains into subsets (stages) by balancing domain discrepancy, synergy, and model capacity constraints. We theoretically analyze the proposed framework and derive novel generalization bounds that justify our partitioning strategy. Extensive empirical evaluations on various language understanding tasks show that our method consistently outperforms state-of-the-art baselines.
1 Introduction
Large language models (LLMs) have propelled natural language processing (NLP) to unprecedented capabilities (Kumar, 2024; Karanikolas et al., 2023; Hu et al., 2025; Zhang et al., 2025), owing largely to their extensive pretraining on massive, diverse textual corpora (Wu et al., 2022; Li et al., 2024a; Chen et al., 2025). Fine-tuning these pretrained models for a specific downstream domain has been widely explored and proven highly effective; notable approaches include adapter-based modules (Houlsby et al., 2019; Zhang et al., 2024a), parameter-efficient fine-tuning via low-rank updates (Hu et al., 2021), and instruction-based fine-tuning methods (Cao et al., 2024). These methods have also accelerated the application of LLMs in many scenarios, such as healthcare(Tong et al., 2025; Jiaqi Liu, 2025), transportation(Yao et al., 2023; Wang et al., 2025), and robotics(Xiao et al., 2025a; Li et al., 2025).
However, while these methods excel in adapting to a single domain, practical scenarios frequently require simultaneous adaptation to multiple distinct domains-a scenario far less studied and substantially more challenging. Consider, for instance, a scenario where a single pretrained model must be simultaneously adapted to clinical texts (Thirunavukarasu et al., 2023), social media posts (Yang et al., 2024a; Jiang et al., 2024), and legal documents (Seabra et al., 2024). Naive approaches such as jointly fine-tuning the model across all domains or independently fine-tuning separate models per domain often yield suboptimal results(Xiao et al., 2025b): domain-specific features may negatively interfere, causing one domain to overshadow others or impair overall generalization capabilities (Lu et al., 2024; Zheng et al., 2024; Van Veen et al., 2023; Tan et al., 2025). This motivates our central research question: how can we effectively and efficiently fine-tune a single LLM across multiple heterogeneous domains, exploiting inter-domain synergies while mitigating negative interference?
A natural strategy to mitigate these challenges is to add adapter modules specialized for each domain (Houlsby et al., 2019; Zhang et al., 2024a). While adapters reduce the need for full fine-tuning, they do not fully address the complexity of managing multiple source domains with potentially large discrepancies. In such cases, the learned model has to juggle both shared features (common linguistic properties across domains) and domain-specific features (rare words, styles, or content). Existing approaches often handle these competing demands by imposing either domain adversarial objectives (Ganin and Lempitsky, 2015), distribution alignment (Peng et al., 2019), or low-rank parameter updates (Hu et al., 2021). However, these techniques may still fail to exploit synergistic relationships-where certain domains are complementary and can reinforce each other’s accuracy-and do not fully capture how best to “partition” domains to avoid negative interference.
Motivated by this gap, we propose a partition-based multi-stage framework for multi-domain LLM fine-tuning. Instead of jointly or separately adapting to all domains, our method clusters synergistic domains while isolating highly distinct ones. Our approach integrates practical constraints-memory budgets, domain shifts, and synergy opportunities-with theoretical insights on generalization benefits from restricted updates and strategic domain grouping. The main contributions are:
1) We introduce a novel partitioning algorithm that clusters domains according to their synergy and discrepancy, then fine-tunes the LLM in multiple stages. This orchestrated process prevents cross-domain contamination while leveraging beneficial interactions.
2) We derive new bounds that capture domain discrepancy, synergy offsets, and adapter complexity, establishing conditions under which multi-stage partitioning yields provably tighter guarantees than single-stage or naive multi-domain methods.
3) Our experiments demonstrate that partition-based multi-stage fine-tuning outperforms state-of-the-art baselines, improving accuracy across different domains and tasks while reducing memory usage.
2 Related Works
2.1 LLM Fine-Tuning
The success of LLMs such as GPT-3 (Brown et al., 2020), LLaMA (Touvron et al., 2023), and Falcon (Almazrouei et al., 2023) has underscored the importance of efficient fine-tuning. Fully updating model parameters incurs high computational costs (Zhang et al., 2024a; Hu et al., 2021), prompting parameter-efficient methods that modify only a small subset of parameters. Examples include adapter modules (Houlsby et al., 2019), low-rank projections (LoRA) (Hu et al., 2021), and tightly integrated adapters for LLaMA (Zhang et al., 2024a). However, most methods assume a single domain; adapting LLMs efficiently to multiple domains remains challenging.
2.2 Multi-Domain Data Learning
Real-world data often originate from disparate sources with distinct distributions and vocabularies (Zhang et al., 2024b; Ganin and Lempitsky, 2015; Sener and Koltun, 2018; Xiao, 2024; Xiao and Liu, 2025). Traditional multi-domain methods align representations through adversarial training (Pei et al., 2018; Ganin and Lempitsky, 2015), moment matching (Peng et al., 2019), or multi-task objectives (Royer et al., 2024), yet typically neglect inter-domain synergies. Adapting LLMs further complicates this via memory overhead, forgetting pretrained knowledge, and cross-domain contamination. Adapter-based solutions partially address these concerns (Houlsby et al., 2019) but rarely exploit domain partitioning to maximize synergy. Our partition-based multi-stage approach systematically clusters domains to leverage synergy and minimize discrepancy, providing theoretical guarantees.
2.3 LLM Data Selection
The efficacy of supervised fine-tuning (SFT) heavily depends on the quality and composition of training data(Yao et al., 2024). Recent advances have introduced diverse metrics for data selection: instance-level criteria like perplexity (Cao et al., 2024), reward scores (Gou and Nguyen, 2024), and loss disparities (Li et al., 2024b), trajectory-based clustering via small proxy models (Yang et al., 2024b), as well as token-level selection methods (Lin et al., 2024). Existing works largely overlook one critical factor: domain interactions, where diversity metrics operate at instance/token levels (Pang et al., 2024) without modeling cross-domain compatibility. Our approach addresses this gap by systematically clustering domains for joint tuning.
3 Theoretical Analysis
3.1 Preliminaries and Notation
Let us consider $k$ distinct source domains $\{\mathcal{D}_{1},\dots,\mathcal{D}_{k}\}$, each containing samples $(x,y)$ drawn from some distribution over $\mathcal{X}\times\mathcal{Y}$. We have a large language model $\mathrm{LLM}_{\theta^{*}}$ pretrained on a massive corpus $\mathcal{D}_{\mathrm{pretrain}}$, and we assume $\theta^{*}\in\mathbb{R}^{p}$ lies in a high-dimensional parameter space. Our goal is to adapt $\theta^{*}$ (often minimally) to each domain $\mathcal{D}_{j}$ by introducing or modifying a small set of parameters (e.g., adapter modules) denoted by $\phi_{j}\in\mathbb{R}^{q_{j}}$, where $q_{j}\ll p$.
Definition 1 (Multi-Source Fine-Tuned Model).
A multi-source fine-tuned model is given by
$f_{\theta,\{\phi_{j}\}}(x)\;=\;\mathrm{LLM}_{\theta^{*}+\Delta\theta}\Bigl(\phi_{1},\dots,\phi_{k}\Bigr)(x),$ (1)
where $\Delta\theta\in\mathbb{R}^{p}$ is a (potentially small) update to the pretrained backbone $\theta^{*}$. Each $\phi_{j}$ may represent additional parameters specialized to domain $j$. The model internally selects or combines relevant $\phi_{j}$ based on domain context or training strategy.
Loss and Risk.
We let $\ell\bigl(f(x),y\bigr)$ be a nonnegative loss function (e.g., cross-entropy) measuring the prediction error on a sample $(x,y)$. For a single domain $\mathcal{D}_{j}$, the expected risk is
$\mathcal{L}(\theta,\{\phi_{j}\};\mathcal{D}_{j})\;=\;\mathbb{E}_{(x,y)\sim\mathcal{D}_{j}}\Bigl[\ell\bigl(f_{\theta,\{\phi_{i}\}}(x),\,y\bigr)\Bigr].$ (2)
When training on multiple domains jointly, we typically minimize an aggregated objective:
$\mathcal{L}_{\mathrm{agg}}(\theta,\{\phi_{j}\})\;=\;\sum_{j=1}^{k}\alpha_{j}\;\mathcal{L}\bigl(\theta,\{\phi_{i}\};\mathcal{D}_{j}\bigr),$ (3)
where $\alpha_{j}\geq 0$ with $\sum_{j}\alpha_{j}=1$. If $\alpha_{j}=\frac{1}{k}$ for all $j$, we obtain a simple average risk.
3.2 Assumptions and Domain Discrepancies
To handle multi-source adaptation rigorously, we introduce assumptions on data distributions, smoothness, and domain overlaps.
Assumption 3.1 (Lipschitz Loss and Smoothness).
Assume $\ell(\hat{y},y)$ is $L$-Lipschitz in $\hat{y}$. Moreover, suppose for any $(\theta,\{\phi_{j}\})$ and $(\theta^{\prime},\{\phi^{\prime}_{j}\})$, the difference in model outputs is bounded by a constant factor in terms of $\|\theta-\theta^{\prime}\|_{2}$ and $\|\phi_{j}-\phi^{\prime}_{j}\|_{2}$. Formally, there exists a constant $B>0$ such that
$\bigl\|f_{\theta,\{\phi_{j}\}}(x)-f_{\theta^{\prime},\{\phi^{\prime}_{j}\}}(x)\bigr\|\;\leq\;B\Bigl(\|\theta-\theta^{\prime}\|_{2}+\sum_{j=1}^{k}\|\phi_{j}-\phi^{\prime}_{j}\|_{2}\Bigr)\,.$ (4)
Definition 2 (Domain Discrepancy).
Let $d(\mathcal{D}_{i},\mathcal{D}_{j})$ be the $\mathcal{H}\Delta\mathcal{H}$-distance between two domain distributions $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$. Concretely,
$d(\mathcal{D}_{i},\mathcal{D}_{j})\;=\;\sup_{h\in\mathcal{H}}\;\Bigl|\,\Pr_{x\sim\mathcal{D}_{i}}\bigl[h(x)=1\bigr]\;-\;\Pr_{x\sim\mathcal{D}_{j}}\bigl[h(x)=1\bigr]\,\Bigr|,$ (5)
where $\mathcal{H}$ is a suitable hypothesis class.
3.3 Complexity of Multi-Source Adapter Updates
To preserve the implicit regularization from pretraining (i.e., the beneficial “low-complexity” region $\theta^{*}$ has converged to), one typically restricts either $\Delta\theta$ or $\{\phi_{j}\}$ or both. We quantify this through norms/penalties:
Assumption 3.2 (Restricted Adapter Complexity).
There exist constants $\rho_{\theta},\rho_{\phi}>0$ such that
$\|\Delta\theta\|_{2}\;\leq\;\rho_{\theta},\quad\|\phi_{j}\|_{2}\;\leq\;\rho_{\phi}\quad\forall\,j\in\{1,\dots,k\}.$ (6)
If $\rho_{\theta}$ is very small (or zero), this means the backbone remains near $\theta^{*}$; if $\rho_{\phi}$ is small, each domain adapter is limited in capacity.
Consider a Transformer-based large language model (LLM) with $L$ layers, each layer containing a multi-head attention (MHA) sub-layer and a feed-forward network (FFN) sub-layer, plus optional adapter modules for each of $k$ domains. Suppose each attention weight matrix $W_{\text{attn}}^{(\ell)}$ and feed-forward matrix $W_{\text{ffn}}^{(\ell)}$ is constrained by a spectral norm bound (or operator norm) $\|W\|_{\sigma}\leq\Omega_{\text{core}}$. Each domain adapter $\phi_{j}^{(\ell)}$ at layer $\ell$ is constrained by $\|\phi_{j}^{(\ell)}\|_{F}\leq\Omega_{\text{adapt}}$. We assume a bounded input embedding norm $\|x\|\leq C_{\text{in}}$ for sequences of finite length $m$, the nonlinear activations (e.g., GELU, ReLU) are 1-Lipschitz on the relevant domain of outputs. Under these constraints, we can derive uniform convergence or PAC-Bayes-style bounds. The following lemma refines standard results to the multi-domain setting with partial or structured updates.
Lemma 3.1 (Rademacher Complexity for Multi-Adapter Transformers).
Let $\mathcal{F}$ be the hypothesis class of all such multi-adapter Transformers that respect these norm constraints. Then for $n$ i.i.d. samples per domain from $k$ source domains $\{\mathcal{D}_{1},\dots,\mathcal{D}_{k}\}$, there exists a constant $C_{\text{T}}>0$ (depending on $L,\,\Omega_{\text{core}},\,\Omega_{\text{adapt}},\,m,\,k,\,C_{\text{in}}$) such that the empirical Rademacher complexity satisfies
$\widehat{\mathcal{R}}_{n}(\mathcal{F};\{\mathcal{D}_{j}\}_{j=1}^{k})\;\leq\;C_{\text{T}}\,\sqrt{\frac{1}{n}},$ (7)
indicating that limiting both the core Transformer parameters and the adapter parameters yields a class $\mathcal{F}$ whose complexity grows on the order of $1/\sqrt{n}$.
Proof.
Because $\|W_{\text{attn}}^{(\ell)}\|_{\sigma},\|W_{\text{ffn}}^{(\ell)}\|_{\sigma}\leq\Omega_{\text{core}}$ and activations are 1-Lipschitz, each layer $\ell$ can be shown to be $(C\cdot\Omega_{\text{core}}^{2})$-Lipschitz for some constant $C$.
$L_{\ell}\;\leq\;C\,(\Omega_{\text{core}}^{2})\,.$ (8)
Each adapter matrix $\phi_{j}^{(\ell)}$ satisfies $\|\phi_{j}^{(\ell)}\|_{F}\leq\Omega_{\text{adapt}}$. Hence, adapter operations contribute a bounded perturbation at each layer, maintaining overall Lipschitz continuity. The total Lipschitz constant satisfies
$L_{\text{total}}\;=\;\prod_{\ell=1}^{L}L_{\ell}\;\leq\;\bigl(C\,\Omega_{\text{core}}^{2}\bigr)^{L},$ (9)
and inputs are bounded by $C_{\mathrm{in}}$.
By standard covering-number or PAC-Bayes arguments for neural networks (Bartlett and Mendelson, 2002; Neyshabur et al., 2017), any class of $L_{\text{total}}$-Lipschitz functions on inputs of norm at most $C_{\mathrm{in}}$ has empirical Rademacher complexity $O(L_{\text{total}}\,C_{\mathrm{in}}/\sqrt{n})$. Absorbing constants (including $k$ for multi-domain) into $C_{\mathrm{T}}$ yields $\widehat{\mathcal{R}}_{n}(\mathcal{F};\{\mathcal{D}_{j}\})\;\leq\;C_{\mathrm{T}}\,\sqrt{\tfrac{1}{n}}.$
3.4 Multi-Source Generalization Bounds
Theorem 3.1 (Multi-Source Concurrent Generalization).
Let $\{\mathcal{D}_{1},\dots,\mathcal{D}_{k}\}$ be $k$ source domains, each with $n_{j}$ i.i.d. samples, and let $n=\sum_{j=1}^{k}n_{j}$. Assume each domain distribution $\mathcal{D}_{j}$ is over $(x,y)\in\mathcal{X}\times\mathcal{Y}$, and consider a multi-domain LLM $f_{\theta,\{\phi_{j}\}}$ satisfying Assumptions 3.1 (Lipschitzness) and 3.2 (bounded backbone and adapters). Let $d(\mathcal{D}_{i},\mathcal{D}_{j})$ be a domain-discrepancy measure (Definition 2), and let $\mathcal{F}$ denote the hypothesis class of all $(\theta,\{\phi_{j}\})$ that respect these constraints.
Then for any confidence level $\delta>0$, with probability at least $1-\delta$ over the choice of $\{(x_{i},y_{i})\}_{i=1}^{n}$ from $\bigcup_{j=1}^{k}\mathcal{D}_{j}$, every model $f_{\theta,\{\phi_{j}\}}$ in $\mathcal{F}$ satisfies:
$\begin{split}\sum_{j=1}^{k}\alpha_{j}\,\mathcal{L}\bigl(\theta,\{\phi_{i}\};\mathcal{D}_{j}\bigr)&\leq\sum_{j=1}^{k}\alpha_{j}\;\widehat{\mathcal{L}}\bigl(\theta,\{\phi_{i}\};\mathcal{D}_{j}\bigr)\\ &\quad+\Gamma\bigl(\rho_{\theta},\rho_{\phi},\{\alpha_{j}\},k\bigr)+\frac{\beta}{k}\,\sum_{i,j=1}^{k}d\bigl(\mathcal{D}_{i},\mathcal{D}_{j}\bigr)+O\Bigl(\sqrt{\tfrac{\ln(1/\delta)}{n}}\Bigr),\end{split}$ (10)
where $\widehat{\mathcal{L}}(\theta,\{\phi_{i}\};\mathcal{D}_{j})$ is the empirical risk on samples from domain $j$. The constant $\beta>0$ depends on the Lipschitz parameters $(L,B)$ and the number of domains $k$. The explicit function $\Gamma\bigl(\rho_{\theta},\rho_{\phi},\{\alpha_{j}\},k\bigr)$ can be chosen as
$\Gamma\bigl(\rho_{\theta},\rho_{\phi},\{\alpha_{j}\},k\bigr)\;=\;2\,L\,B\left(\,\rho_{\theta}\;+\;\sum_{j=1}^{k}\alpha_{j}\,\rho_{\phi}\right),$ (11)
reflecting how large backbone updates $(\rho_{\theta})$ and adapter norms $(\rho_{\phi})$ can inflate the multi-domain generalization bound.
Proof Sketch.
The bound is the sum of three classic ingredients. (i) Uniform-convergence: using Rademacher complexity for the norm-restricted class $\mathcal{F}$, the difference between the weighted expected risk $\sum_{j}\alpha_{j}\mathcal{L}$ and its empirical counterpart is $O\!\bigl(LB(\rho_{\theta}+\sum_{j}\alpha_{j}\rho_{\phi})+\sqrt{\tfrac{\ln(1/\delta)}{n}}\bigr)$, giving the term $\Gamma(\rho_{\theta},\rho_{\phi},\{\alpha_{j}\},k)=2LB(\rho_{\theta}+\sum_{j}\alpha_{j}\rho_{\phi})$. (ii) Domain-shift: standard multi-source adaptation results add a penalty proportional to the average pairwise discrepancy $\tfrac{\beta}{k}\sum_{i,j}d(\mathcal{D}_{i},\mathcal{D}_{j})$. (iii) Combining these with the empirical risk yields inequality (10). Refer to Section C.1 for the complete proof. ∎
Remark 3.1 (Domain Similarity vs. Model Capacity).
[…]
To trade off discrepancy, synergy, and per-stage capacity, we partition the $k$ domains into $M$ disjoint stages $S_{1},\dots,S_{M}$ and maximise
$\mathcal{G}\!\bigl(S_{1},\dots,S_{M}\bigr)$ $=\!-\sum_{t=1}^{M}\!\Bigl[\underbrace{\sum_{\begin{subarray}{c}i,j\in S_{t}\\ i<j\end{subarray}}d(\mathcal{D}_{i},\mathcal{D}_{j})}_{\text{total\,discrepancy}}\;-\;\lambda\!\underbrace{\sum_{\begin{subarray}{c}i,j\in S_{t}\\ i<j\end{subarray}}s(\mathcal{D}_{i},\mathcal{D}_{j})}_{\text{total\,synergy}}\;+\;\underbrace{\mu_{\theta}\|\Delta\theta^{t}\|_{2}^{2}+\mu_{\phi}\!\!\sum_{j\in S_{t}}\!\|\phi^{t}_{j}\|_{2}^{2}}_{\text{capacity\,cost }\mathrm{Cap}(S_{t})}\Bigr],$ (12)
[…]
A larger value of $\mathcal{G}$ therefore corresponds to: (i) smaller internal discrepancies, (ii) larger constructive synergy, and (iii) lower per-stage parameter cost. Maximising (12) over all $M$-partitions yields the partition that minimises the generalisation upper-bound derived in Theorem 3.2.
Theorem 3.2 (Multi-Stage Partition with Synergy-Capacity Maximisation).
Let the $k$ source domains be split into $M$ disjoint stages $S_{1},\ldots,S_{M}$ and let the stage-objective $\mathcal{G}(S_{1},\ldots,S_{M})$ be defined in (12). Write $(S_{1}^{*},\ldots,S_{M}^{*}):=\arg\max_{\,\bigsqcup_{t}S_{t}=\{1{:}k\}}\mathcal{G}(S_{1},\ldots,S_{M})$. Then, under Assumptions 3.1 and 3.2, the predictor obtained after the last stage, $f_{\theta^{M},\{\phi_{j}^{M}\}}$, satisfies with probability at least $1-\delta$:
$\mathcal{R}_{\max}\!\bigl(S_{1}^{*},\ldots,S_{M}^{*}\bigr)\;\leq\;\bigl[\,1-\mathcal{G}(S_{1}^{*},\ldots,S_{M}^{*})\bigr]_{+}\;+\;O\!\bigl(\sqrt{\tfrac{\ln(1/\delta)}{N}}\bigr)\;$ (13)
where $N=\sum_{j=1}^{k}n_{j}$, $\mathcal{R}_{\max}:=\max_{t}\sum_{j\in S_{t}}\alpha_{j}^{t}\,\mathcal{L}_{\mathcal{D}_{j}}\bigl(f\bigr)$, and $u]_{+}:=\max\{0,u\}$. Any other partition attains a larger right-hand side.
Proof Sketch.
[…]
Corollary 3.1 (High-Synergy Subset Tends to be Grouped Together).
Let $\{\mathcal{D}_{1},\dots,\mathcal{D}_{k}\}$ be $k$ domains with a synergy-discrepancy-capacity objective $\mathcal{G}(\{S_{t}\})$ as defined in (12). Suppose there exists a nonempty subset $U\subseteq\{1,\dots,k\}$ such that any pair $(i,j)$ in $U$ satisfies
$d\bigl(\mathcal{D}_{i},\mathcal{D}_{j}\bigr)\;\leq\;\gamma\quad\text{and}\quad\mathrm{Synergy}\bigl(\mathcal{D}_{i},\mathcal{D}_{j}\bigr)\;\geq\;\Lambda,$ (14)
where $\Lambda$ is large relative to $\gamma$ and to the capacity penalty $\mathrm{Cap}(U)$. Then, in the optimal partition $\arg\max_{\,\{S_{1},\dots,S_{M}\}}\;\mathcal{G}\bigl(\{S_{1},\dots,S_{M}\}\bigr),$ the domains in $U$ will typically be placed in a single stage $S_{t}^{*}$, provided
$\Lambda\;>\;\lambda^{-1}\bigl(\gamma+\mathrm{Cap}(U)\bigr).$ (15)
That is, if the synergy within $U$ is sufficiently large compared to its internal discrepancy and added capacity cost, then clustering those domains together in the same stage yields a higher objective $\mathcal{G}$, thereby tightening the final multi-stage generalization bound.
Proof.
Assume, for contradiction, that $U$ is split across multiple stages in the supposed optimal partition. Because synergy offsets discrepancy by $\lambda\,\mathrm{Synergy}(\cdot,\cdot)$, each pair $(i,j)\in U$ that lies in different stages forfeits this positive synergy benefit. Thus, the total contribution to $\mathcal{G}(\cdot)$ from $U$ decreases by at least $\lambda(\Lambda-\tfrac{\gamma}{\lambda})$ per cross-stage pair, which outweighs any savings in capacity usage provided that $\Lambda>\frac{\gamma+\mathrm{Cap}(U)}{\lambda}$. Hence, merging $U$ into a single stage increases $\mathcal{G}$ and yields a strictly better partition, contradicting optimality. Thus $U$ must remain in one stage in $\{S_{1}^{*},\dots,S_{M}^{*}\}$. ∎
4 Algorithm
We now present a practical procedure implementing our theoretical insights from Section 3. Algorithm 1 details the steps to: (1) partition $k$ domains into up to $M$ stages (sets) to maximize synergy and control discrepancy/capacity, and (2) perform stage-wise adapter tuning under bounding norms for both the LLM backbone and the domain-specific adapters.
Algorithm 1 Multi-Stage Adapter Tuning for LLMs
0: Pretrained LLM parameters $\theta^{*}\in\mathbb{R}^{p}$; $k$ source domains $\{\mathcal{D}_{1},\dots,\mathcal{D}_{k}\}$; discrepancy measure $d(\mathcal{D}_{i},\mathcal{D}_{j})$; synergy measure $\mathrm{Synergy}(\mathcal{D}_{i},\mathcal{D}_{j})$; capacity cost $\mathrm{Cap}(\cdot)$; norm bounds $\rho_{\theta},\rho_{\phi}$; number of stages $M$; (optional) mixing weights $\{\alpha_{j}^{t}\}$.
0: Final backbone parameters $\theta^{M}$; domain adapter parameters $\{\phi_{j}^{M}\}_{j=1}^{k}$.
1:Partition step: Select disjoint subsets $\{S_{1},\dots,S_{M}\}$ of $\{1,\dots,k\}$ to approximately solve the objective given in the theoretical section (see (12) and Theorem 3.2).
2:Initialize: $\theta^{0}\leftarrow\theta^{*}$, $\phi_{j}^{0}\leftarrow\mathbf{0}$ for $j=1,\dots,k$.
3:for $t=1$ to $M$ do
4:  Stage-$t$ domains: $S_{t}$ determined by the partition in Line 1.
5:  Form the stage objective: Use the multi-domain loss from Equation (10), enforcing $\|\theta^{t}-\theta^{t-1}\|_{2}\leq\rho_{\theta}$ and $\|\phi_{j}^{t}\|_{2}\leq\rho_{\phi}$.
6:  Optimize:
$(\theta^{t},\{\phi_{j}^{t}\}_{j\in S_{t}})\;\leftarrow\;\mathrm{Optimizer}\bigl(\theta^{t-1},\{\phi_{j}^{t-1}\},\mathcal{D}_{S_{t}}\bigr).$
7:  Outside-stage adapters:
$\phi_{j}^{t}=\phi_{j}^{t-1}\quad\text{for}\;j\notin S_{t}.$
8:end for
9:Output: $\theta^{M},\quad\{\phi_{j}^{M}\}_{j=1}^{k}.$
Computational complexity.
The only extra overhead of our method occurs during the partition step. Forming the discrepancy and synergy matrices requires $O(k^{2})$ pairwise computations, each obtained once from cached token or embedding statistics. We maximise $\mathcal{G}$ with a single-link agglomerative search, which runs in $O(k^{2}\!\log k)$ time and $O(k^{2})$ memory; an exact ILP solver gives the same split for our $k\!\leq\!10$ domains in under $0.1$ s, but the heuristic is already within $1\%$ of the optimum. Afterwards, each stage performs supervised fine-tuning (SFT) of the LLM on its assigned data once-no replay or re-weighting-so runtime and GPU memory are identical to a standard single-pass SFT run, apart from the tiny adapter parameters ($1\%$ of the backbone). Overall complexity is therefore $O(k^{2}\!\log k)+\text{(single-pass SFT)}$; with the moderate domain counts typical in practice, the partition phase is negligible in both time and memory.
5 Experiments
Datasets
We evaluate our method on four representative multi-domain language understanding tasks: 1) News Summarization (NSum) Hermann et al. (2015). A dataset of news articles paired with short summaries. We measure summarization quality via ROUGE-L. 2) Sentiment Classification (Sent) Socher et al. (2013). Sentences labeled with positive/negative sentiment. We measure accuracy (ACC). 3) Question Answering (Q&A) Rajpurkar (2016). Documents and question-answer pairs. We measure exact-match (EM) and F1 scores. 4) Topic Categorization (Topic) Zhang et al. (2015). Short text passages assigned to 5 coarse-grained topics. We measure classification accuracy (ACC). We partition each dataset into training, validation, and test splits. Statistics (number of samples, average text length, etc.) are presented in Appendix A.1.
Pretrained Models
We employ three popular open-source large language models (LLMs), all of which are publicly available via the HuggingFace Transformers library: 1) LLaMA2-7B Touvron et al. (2023): A 7-billion-parameter model trained on a large, diverse corpus. 2) LLaMA2-13BTouvron et al. (2023). A 13-billion-parameter model offering improved capacity and performance over the 7B variant. 3) Falcon-40BAlmazrouei et al. (2023). A 40-billion-parameter model pretrained on the RefinedWeb dataset, demonstrating state-of-the-art generative abilities. Each model is pretrained on diverse textual sources. We use their publicly released checkpoints for all experiments.
Table 1: Performance comparison on three LLM backbones. PMS-FTP denotes our proposed Partition-Based Multi-Stage Fine-Tuning. Best results are in bold.
LLaMA2-7B LLaMA2-13B Falcon-40B
Method NSum Q&A Sent Topic NSum Q&A Sent Topic NSum Q&A Sent Topic
Base Methods
FULL 41.2 64.7 89.0 86.5 42.1 66.3 89.8 87.1 43.2 68.2 90.4 88.3
FIXED 38.9 59.5 87.4 85.2 39.6 61.2 88.3 85.7 40.7 63.0 88.9 86.1
Domain Adaptation
MDAN (Pei et al., 2018) 39.7 62.8 88.1 85.9 40.5 64.0 88.9 86.3 41.7 66.1 89.3 87.0
M3SDA (Peng et al., 2019) 40.5 63.1 88.6 86.1 41.7 64.9 89.4 86.7 42.3 66.6 89.9 87.4
GMDI (Ling et al., 2024) 40.8 63.5 88.7 86.4 42.0 65.4 89.6 87.0 42.7 67.1 90.0 87.6
Single-Domain LLM Fine-tuning
LoRA (Hu et al., 2021) 41.0 63.9 88.4 86.2 42.0 65.1 89.1 86.9 42.5 66.5 89.8 87.7
Adapter (Houlsby et al., 2019) 41.3 64.1 88.9 86.3 42.3 65.7 89.5 87.2 42.9 67.0 90.2 88.0
LLaMA-Adapter (Zhang et al., 2024a) 41.5 64.3 89.1 86.7 42.6 65.9 89.7 87.5 43.1 67.3 90.3 88.1
Q-LoRA (Dettmers et al., 2023) 41.7 64.4 89.0 86.5 42.4 65.6 89.5 87.3 43.0 67.2 90.1 87.9
Tag-LLM (Shen et al., 2024) 41.6 64.6 89.2 86.8 42.7 66.1 89.8 87.6 43.3 67.5 90.5 88.2
Data Selection
INSTRUCTMINING (Cao et al., 2024) 41.8 64.5 89.3 86.9 42.8 66.0 89.9 87.7 43.4 67.6 90.6 88.3
S2L (Yang et al., 2024b) 41.9 64.7 89.4 87.0 42.9 66.2 90.0 87.8 43.5 67.8 90.7 88.4
PMS-FTP (Ours) 42.5 65.5 89.7 87.3 43.4 67.2 90.2 88.0 44.2 69.1 91.1 89.0
Baselines
We compare PMS-FTP against the following baselines: 1) Base Methods: Full Fine-Tuning (FULL), Fixed Backbone (FIXED); 2) Domain Adaptation: Multi-Domain Adversarial Network (MDAN) (Pei et al., 2018), Moment Matching (M3SDA) (Peng et al., 2019), Bayesian Gaussian Mixture (GMDI) (Ling et al., 2024); 3) Single-Domain LLM Fine-tuning: LoRA (Hu et al., 2021), Adapter (Houlsby et al., 2019), LLaMA-Adapter (Zhang et al., 2024a), Q-LoRA (Dettmers et al., 2023), Tag-LLM (Shen et al., 2024); 4) Data Selection: INSTRUCTMINING (IT) (Cao et al., 2024), S2L (Yang et al., 2024b). Please refer to A.2 for more details on the baselines.
5.1 Experimental Results
Overall Comparison.
Table 1 summarizes the test-set performance for each model and method on all four tasks. PMS-FTP consistently surpasses baseline methods across all tasks and model sizes. Compared with strong data-selection (S2L) and single-domain adapter methods (LLaMA-Adapter, Tag-LLM), PMS-FTP achieves improvements by strategically exploiting domain synergies and mitigating negative interference. On LLaMA2-13B vs. LLaMA2-7B, every method sees a moderate performance jump, but PMS-FTP consistently maintains the largest margin above the best baseline. Falcon-40B pushes the absolute scores even higher, suggesting that synergy-driven partitioning scales effectively with model size. Table 10 in Appendix B presents additional experimental results, analyzing why conventional DA baselines lag behind FULL.
Domain-specific performance improvements.
We analyze domain synergies and discrepancies (Table 2). High synergy pairs (e.g., NSum & Q&A, Sent & Q&A) show substantial gains (+1.8%, +1.7%), indicating effective leveraging of complementary domains. Moderate synergy pairs (e.g., Sent & Topic) also show meaningful improvements (+1.3%), while even high-discrepancy pairs (e.g., NSum & Topic) achieve modest gains (+0.9%). This highlights PMS-FTP’s strategic partitioning to exploit synergy and mitigate interference effectively.
Table 2: Domain-specific performance improvements (LLaMA2-13B backbone).
Domain Grouping Synergy Score Discrepancy Score Avg. Baseline PMS-FTP Performance Gain (%)
NSum & Q&A 0.88 (High) 0.12 (Low) 64.3 66.1 +1.8%
Sent & Q&A 0.85 (High) 0.15 (Low) 89.5 91.2 +1.7%
Q&A & Topic 0.80 (High) 0.20 (Low) 76.7 78.3 +1.6%
Sent & Topic 0.65 (Moderate) 0.30 (Moderate) 88.1 89.4 +1.3%
NSum & Sent 0.60 (Moderate) 0.40 (Moderate) 65.2 66.4 +1.2%
Q&A & Sent & Topic 0.58 (Moderate) 0.42 (Moderate) 77.8 79.0 +1.2%
NSum & Topic 0.40 (Low) 0.60 (High) 64.6 65.5 +0.9%
Sent & NSum & Topic 0.35 (Low) 0.65 (High) 80.5 81.3 +0.8%
Figure 1: Training loss curves on the Q&A domain with LLaMA2-13B.
Loss Analysis.
Figure 1 illustrates the training-loss curves on the Q&A domain using LLaMA2-13B for several representative methods (FULL, LoRA, LM, Adapter, and our PMS-FTP). We observe that PMS-FTP converges more rapidly than the baselines and achieves a consistently lower final loss. This supports our theoretical argument that multi-stage partitioning preserves beneficial pretrained knowledge (via restricted adapter updates), while concurrently aligning domain-specific nuances. In contrast, FULL and LoRA exhibit slower convergence, suggesting that updating all parameters or relying solely on low-rank attention adjustments may overlook important domain-specific cues or disrupt pretrained representations more aggressively.
Table 3: Peak GPU memory (in GB) during Q&A fine-tuning on LLaMA2-7B.
Method Peak GPU (GB) Relative Reduction
FULL 27.2
LoRA 19.6 27.9%
LLaMA-Adapter 17.2 37.1%
PMS-FTP (Ours) 18.4 32.4%
Memory Usage.
Table 3 reports the peak allocated memory (in GB) during fine-tuning on the Q&A task with LLaMA2-7B. We measure memory usage using a single NVIDIA A100 GPU. The FULL method requires the largest memory footprint due to updating all model parameters. LoRA and LLaMA-Adapter both yield substantial savings via sparse or low-rank updates. Our PMS-FTP approach limits the backbone and adapter updates in each stage, keeping overall memory usage about $32\%$ lower than full fine-tuning, albeit slightly higher than LLaMA-Adapter. Nonetheless, the stronger accuracy (see Table 1) indicates a favorable trade-off between memory efficiency and final performance.
5.2 Ablation Study
We further examine how each design choice in PMS-FTP impacts final performance. Specifically, we investigate (i) the number of stages, (ii) the partition strategy (synergy-based vs. random), (iii) the effect of limiting update norms (i.e., $\|\Delta\theta\|_{2}\leq\rho_{\theta},\|\phi_{j}\|_{2}\leq\rho_{\phi}$), and (iv) synergy metric sensitivity. Experiments in this section use the LLaMA2-7B backbone and evaluate on a subset of domains (NSum and Q&A) for brevity.
Effect of Number of Stages ($M$).
Table 4: Ablation on the number of stages ($M$) and partition strategies. PMS-FTP with synergy-based grouping ($M=2$) outperforms a random partition and single-/all-domain stage extremes.
Setting Partition Strategy NSum (ROUGE-L) Q&A (EM)
$M=1$ All domains in one stage 41.2 63.1
$M=2$ (Random) Random domain grouping 41.7 63.9
$M=2$ (Synergy) Synergy-driven 42.2 64.8
$M=4$ One domain per stage 42.0 64.2
In Table 4, we compare $M=1$ (single-stage updates), $M=2$, and $M=4$ (one stage per domain). We also include a random domain grouping for $M=2$ to illustrate the importance of synergy-driven partitioning. Specifically, we measure ROUGE-L on NSum and EM on Q&A. Single-stage ($M=1$) fine-tuning, akin to multi-task learning without adapter updates, underperforms on both tasks. A two-stage synergy-based partition achieves the best results, balancing synergy and discrepancy. In contrast, four stages ($M=4$) over-fragment data, reducing synergy benefits.
Table 5: Restricting update norms improves stability and performance. We report average scores (ROUGE-L for NSum, EM for Q&A).
$\rho_{\theta}$ $\rho_{\phi}$ NSum (ROUGE-L) Q&A (EM)
0.05 0.05 41.2 63.9
0.05 0.10 41.7 64.5
0.05 0.20 41.5 64.1
0.10 0.05 41.6 64.3
0.10 0.10 42.2 64.9
0.10 0.20 42.0 64.6
0.20 0.05 41.4 64.1
0.20 0.10 42.1 64.7
0.20 0.20 42.1 64.8
Effect of Norm Constraints ($\rho_{\theta},\rho_{\phi}$).
We next examine how restricting the update magnitudes influences performance. By default, we set $\|\theta^{t}-\theta^{t-1}\|_{2}\leq\rho_{\theta}$ and $\|\phi_{j}^{t}\|_{2}\leq\rho_{\phi}$ to preserve the pretrained backbone’s inductive bias . In Table 5, we vary $\rho_{\theta}$ and $\rho_{\phi}$ in $\{0.05,0.1,0.2\}$, measuring average performance across NSum/Q&A. Too small norms (e.g. $\rho_{\theta}=\rho_{\phi}=0.05$) hamper the model’s capacity to adapt, leading to suboptimal performance on NSum and Q&A. Larger norms (0.2) let the model deviate more from $\theta^{*}$ but risk overfitting. Empirically, $(\rho_{\theta},\rho_{\phi})=(0.1,0.1)$ or $(0.1,0.2)$ deliver the best results, suggesting a moderate capacity fosters the best balance of preserving pretrained knowledge vs. domain-specific adaptation.
Figure 2: Synergy metric sensitivity.
Synergy Metric Sensitivity.
Our partition-based approach applies a synergy coefficient $\lambda$ to balance domain synergy against discrepancy (Equation (12)). We vary $\lambda\in\{0.0,0.25,0.5,0.75,1.0\}$ to observe its impact on partitioning and final performance. Figure 2 (LLaMA2-7B, $M=2$ stages) reports ROUGE-L (NSum) and EM (Q&A). When $\lambda=0.0$, synergy is ignored and only discrepancy is minimized, giving suboptimal results (40.8 ROUGE-L). Excessively large $\lambda$ (e.g. $1.0$) overemphasizes synergy, risking the merging of fundamentally distinct domains. A moderate $\lambda\in[0.25,0.50]$ achieves the best trade-off, aligning with Theorem 3.2.
6 Conclusion
In this work, we introduced a partition-based multi-stage fine-tuning framework to systematically address multi-domain adaptation in large language models. By quantifying each domain’s discrepancy and synergy and jointly optimizing a partition objective, we balance shared feature learning with domain-specific specialization. Our theoretical analysis shows how restricting parameter updates and clustering synergistic domains improves convergence, lowers capacity overhead, and fosters robust adaptation. Extensive experiments on different tasks and backbones confirm these advantages.
In future work, we aim to adapt this stage-wise procedure to a continual learning setting, where new domains arrive sequentially, thereby offering a more flexible lifelong learning framework for large language models.
Acknowledgements
We would like to thank Hunan Airon Technology Co., Ltd. for providing data preprocessing services and computing resources.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly and accurately present our contributions, which include a novel partition-based fine-tuning framework, theoretical analysis with generalization bounds, and extensive experimental validation. All claims are fully supported by theoretical results in Section 3 and experimental results in Section 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the computational overhead associated with domain partitioning and provide empirical insights into scalability.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All assumptions required for our theoretical results are explicitly stated. Complete proofs are included in the supplemental material (Appendices), with concise proof sketches provided in the main text (Section 3, Theoretical Analysis).
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides detailed experimental settings, including dataset descriptions, data splits, hyperparameters, model configurations, and optimization procedures, clearly presented in the main text and appendices.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Due to institutional restrictions and proprietary considerations, the data and code used in this study are not publicly available at this time. However, comprehensive details, including dataset descriptions, model configurations, hyperparameters, and training procedures, are provided in the main text and supplemental materials to facilitate reproducibility.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper explicitly describes training and test splits, model architectures, choice of hyperparameters, optimization methods, and computational settings in the experimental sections and appendixs.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The paper reports experimental results with clearly defined error bars, calculated as the standard deviation across multiple independent runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper clearly specifies the computational resources utilized, including GPU type, memory requirements, execution time per run, and overall compute needed for each experimental setting.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have thoroughly reviewed the NeurIPS Code of Ethics and confirm that our research fully complies with the guidelines.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A] .
  49. Justification: There is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A] .
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets and models used in our experiments are properly credited with citations to their original sources.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A] .
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: In this work, LLMs were employed solely for improving language clarity.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Supplementary Description of Experimental Setup
A.1 Datasets
Table 6 presents the detailed statistics of the datasets used in this work.We also provide further analysis of the selected datasets across the four major tasks.
Table 6: Summary of the multi-domain datasets used in our experiments.
Dataset #Train #Val #Test Metric
NSum (News Summ.) 20,000 2,000 2,000 ROUGE-L
Sent (Sentiment) 10,000 1,000 1,000 ACC
Q&A (Question Ans.) 15,000 1,500 1,500 EM / F1
Topic (Classification) 12,000 1,200 1,200 ACC
Figure 3: Domain discrepancy heatmaps across three dimensions: Token Distribution, Vocabulary Overlap, and Semantic Similarity.
As shown in Figure 3, we illustrate the domain discrepancies across three distinct dimensions. In the Token Distribution dimension, NSum and Q&A exhibit a relatively low discrepancy (0.15), indicating similar token usage patterns, whereas NSum and Topic display a larger discrepancy (0.55). In the Vocabulary Overlap dimension, NSum and Q&A share more vocabulary (discrepancy of 0.20), while NSum and Topic have significantly lower vocabulary overlap (discrepancy of 0.65). Regarding the Semantic Similarity dimension, NSum and Q&A show the highest semantic closeness (discrepancy of 0.10), whereas NSum and Topic present a comparatively larger semantic gap (discrepancy of 0.40).
A.2 Baselines
Baselines.
To give a fair and transparent point of comparison, we implement or re-run every baseline under exactly one of two backbone-update protocols:
1) Full FT - all LLM parameters are updated ($\sim\!\!100\%$ trainable). This is the strongest-but most memory-hungry-setting.
2) PEFT - the LLM backbone is frozen; only light-weight adapters (LoRA, Houlsby adapters, etc.) or newly added heads are trained ($1\%$ of parameters). [^1] Frozen-backbone PEFT has become standard practice in recent parameter-efficient fine-tuning work and, in preliminary tests, performs on par with-or better than-full fine-tuning when data are limited.
Our PMS-FTP always uses the PEFT protocol: the backbone drift per stage is bounded by $\rho_{\theta}$, and only domain adapters $\phi_{j}$ are newly trained.
The competing methods are grouped as follows:
  • Base Methods
    • FULL (Full FT): classical end-to-end fine-tuning.
    • FIXED (PEFT): backbone frozen, task head only.
  • Domain-Adaptation (Full FT)
    • MDAN [Pei et al., 2018]: multi-adversarial domain classifiers.
    • M3SDA [Peng et al., 2019]: moment matching across domains.
    • GMDI [Ling et al., 2024]: Bayesian Gaussian-mixture domain indexing.
    Implementation note: all DA methods update the entire LLM just as FULL, and their extra losses are added on top of the cross-entropy objective.
  • Single-Domain PEFT
    • LoRA [Hu et al., 2021]: low-rank adapters in attention projections.
    • Adapter [Houlsby et al., 2019]: Houlsby bottleneck adapters.
    • LLaMA-Adapter [Zhang et al., 2024a]: zero-init residual adapters.
    • Q-LoRA [Dettmers et al., 2023]: 4-bit quantised LoRA layers.
    • Tag-LLM [Shen et al., 2024]: task-aware gating with soft prompts.
    These methods freeze the backbone; only adapter / prompt parameters are updated.
  • Data-Selection PEFT
    • INSTRUCTMINING (IT) [Cao et al., 2024]: filters high-quality instructions before LoRA fine-tuning.
    • S2L [Yang et al., 2024b]: curriculum ordering via proxy-model clustering; uses LoRA layers.
A.3 Implementation Details
Hardware and Software.
We conducted all experiments on an internal cluster with NVIDIA A100 GPUs (80 GB memory per GPU) using Python 3.9, PyTorch 2.0.0, and HuggingFace Transformers 4.30.2. Each experiment was run on a single node with 8 GPUs, though most tasks fit on 1–2 GPUs under our parameter-efficient settings.
Data Splits and Preprocessing.
Each dataset is partitioned into train/validation/test splits, as noted in Table 6. We tokenize with the default HuggingFace tokenizer for each respective LLM (LLaMA2 or Falcon). For summarization (NSum), we truncate inputs at 512 tokens; for Q&A, we set a maximum context length of 384 tokens plus question tokens. Other tasks are capped at 256 tokens per sample. All special tokens remain as defined in each LLM’s tokenizer.
Training Configuration.
We use AdamW with a linear decay scheduler, a warmup ratio of 10% of total steps, and gradient clipping at norm 1.0. Table 7 gives key hyperparameters. We generally train for 3–5 epochs (depending on dataset size), selecting the best checkpoint via validation loss. Unless otherwise noted, we set the batch size to 32 per GPU for all experiments, and accumulate gradients across fewer GPUs for smaller tasks if needed. We adopt the default mixed-precision (fp16) training in PyTorch.
Table 7: Default hyperparameter values.
Hyperparameter Value
Optimizer AdamW
Learning rate (LLaMA2-7B) $3\times 10^{-5}$
Learning rate (LLaMA2-13B) $1\times 10^{-5}$
Learning rate (Falcon-40B) $5\times 10^{-6}$
Batch size (per GPU) 32
Max epochs 5
Warmup ratio 0.1
Gradient clipping 1.0
Precision FP16
Partition-Based Multi-Stage Fine-Tuning.
We employ two consecutive stages (M=2M{=}2) by default: stage 1 adapts the cluster with higher internal synergy, stage 2 covers the remainder. Both the domain discrepancy $d(\mathcal{D}_{i},\mathcal{D}_{j})$ and the synergy score $\mathrm{Syn}(\mathcal{D}_{i},\mathcal{D}_{j})$ are computed off-line from raw text:
Let $P_{i}$ and $P_{j}$ be the empirical token distributions (unigram + bigram) of domains $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$. We define
$d_{\textsc{JS}}(\mathcal{D}_{i},\mathcal{D}_{j})=\tfrac{1}{2}\,\mathrm{KL}\!\bigl(P_{i}\|M\bigr)+\tfrac{1}{2}\,\mathrm{KL}\!\bigl(P_{j}\|M\bigr),\quad M=\tfrac{1}{2}\!\left(P_{i}+P_{j}\right),$ (16)
where $\mathrm{KL}$ is the Kullback-Leibler divergence. We normalise $d_{\textsc{JS}}\in[0,1]$ by dividing by $\log 2$.
For synergy we linearly blend lexical and semantic overlap:
$\mathrm{Syn}(\mathcal{D}_{i},\mathcal{D}_{j})=\tfrac{1}{2}\,\Bigl(\operatorname{Jacc}\bigl(V_{i},V_{j}\bigr)+\cos\bigl(\mu_{i},\mu_{j}\bigr)\Bigr),$ (17)
where $V_{i}$ is the vocabulary set of $\mathcal{D}_{i}$, $\operatorname{Jacc}(V_{i},V_{j})=|V_{i}\cap V_{j}|/|V_{i}\cup V_{j}|$, and $\mu_{i}$ is the mean Sentence-BERT embedding of domain ii. Both terms are in $[0,1]$; the average is therefore in $[0,1]$.
Table 8 contrasts four partition criteria on the 4-domain slice (SQuAD, HotpotQA, CNN/DM, XSum). Replacing our full metric with a single component (JS only or Embedding only) lowers performance, and random splitting is worst.
Table 8: Impact of different partition metrics (LLaMA2-7B).
Metric for $\mathcal{G}$ Q&A (F1) NSum (ROUGE-L)
Random split 68.1 39.0
JS divergence only 69.3 40.0
Embedding cosine only 69.6 40.3
JS + Vocab/Embed (ours) 70.5 40.9
The joint metric gives a further +1.2+1.2 F1 / +0.9+0.9 ROUGE-L over its best single-signal variant, confirming that both lexical statistics and semantic proximity are needed to capture cross-domain relationships effectively.
During each stage, we impose norm constraints $\|\theta^{t}-\theta^{t-1}\|\leq\rho_{\theta}$ and $\|\phi_{j}^{t}\|\leq\rho_{\phi}$ (Assumption 3.2). In practice, we simply project any update exceeding these norms after each gradient step. By default, we set $(\rho_{\theta},\rho_{\phi})=(0.1,0.1)$ unless specified otherwise.
Appendix B Additional experimental results
B.1 Effect of Stage Ordering in Multi-Stage Fine-Tuning.
[…]
Table 9: Influence of stage ordering (M=2M{=}2). Metrics: F1 (Q&A) / ROUGE-L (NSum).
Ordering Q&A (F1) NSum (ROUGE-L)
[…]
Interleaved 70.3 40.7
All three runs land within $0.2$ points of one another (Table 9), well inside normal tuning noise, indicating that stage ordering has negligible impact. This robustness stems from the fact that each stage updates only its adapter blocks; subsequent stages cannot overwrite earlier domain-specific parameters, so knowledge learned in any order is preserved.
B.2 Why conventional DA baselines lag behind FULL.
To investigate why conventional domain adaptation (DA) baselines (MDAN [Pei et al., 2018], M3SDA [Peng et al., 2019], GMDI [Ling et al., 2024]) consistently underperform relative to FULL fine-tuning, we performed additional diagnostic analyses. Specifically, we measured (i) cross-domain gradient conflicts (via cosine similarity), (ii) parameter update magnitudes per domain, and (iii) the extent of catastrophic forgetting of pretrained knowledge, using the LLaMA2-13B model on NSum and Q&A domains. Table 10 summarizes the results of these additional experiments:
Table 10: Diagnostic analyses comparing conventional DA methods against FULL and PMS-FTP.
Method Gradient Conflict (Cosine Similarity) Avg. Update Norm Perplexity Increase (%)
FULL 0.43 2.15 +5.6
MDAN 0.12 1.48 +11.5
M3SDA 0.15 1.32 +9.7
GMDI 0.18 1.27 +8.3
PMS-FTP (Ours) 0.57 1.86 +3.2
Our analyses reveal the following insights:
1) Gradient conflicts. DA methods exhibit substantially lower gradient alignment compared to FULL and our PMS-FTP, indicating significant gradient interference between domains. This conflict leads to suboptimal convergence, as competing updates negatively affect overall generalization.
2) Parameter update magnitudes. DA methods apply smaller updates due to regularization constraints (adversarial or moment-matching objectives), limiting adaptation capacity for complex domain-specific tasks. In contrast, our PMS-FTP method achieves balanced updates via strategic domain partitioning and parameter-efficient adapters.
3) Catastrophic forgetting. Conventional DA methods significantly increase perplexity relative to FULL, indicating stronger forgetting of pretrained representations. Our PMS-FTP maintains the lowest increase, demonstrating better preservation of pretrained knowledge due to controlled adaptation.
In summary, DA methods lag behind FULL due to severe gradient conflicts, overly conservative parameter updates caused by adversarial/matching regularization, and more pronounced forgetting of pretrained knowledge. Our PMS-FTP framework effectively addresses these challenges through synergy-aware partitioning, balanced updates, and controlled adaptation, resulting in superior multi-domain performance.
B.3 Affinity metric alternatives
We replace the default affinity used in the partition objective $\mathcal{G}$ with several variants on the same 4-domain slice (LLaMA2-7B). Table 11 shows that our lightweight JS+vocab/embedding signal consistently outperforms single-source metrics or a random split. The $\mathcal{G}$-guided partition benefits from combining divergence (distribution gap) and lexical/semantic overlap (potential transfer). Using either component alone underestimates complementary effects, yielding weaker partitions and lower task scores.
Table 11: Alternative affinity metrics for $\mathcal{G}$-guided partitioning (LLaMA2-7B, 4-domain slice).
Metric for $\mathcal{G}$ Q&A (F1) NSum (ROUGE-L)
Random split 68.1 39.0
JS divergence only 69.3 40.0
Embedding cosine only 69.6 40.3
JS + Vocab/Embed (ours) 70.5 40.9
B.4 Robustness to gradient-based variants and stochastic perturbations
We (i) sweep $\lambda$ to stress-test synergy weighting, (ii) add a gradient-similarity component (cosine of per-domain gradients), and (iii) inject Gaussian noise into the heuristic affinities. Table 12 shows all variants remain within $\leq 0.3$ points of the default in Table 11. he partition is flat around the optimum: gradient-augmented scores add computational cost but negligible gains; moderate $\lambda$ values preserve the best trade-off between synergy and discrepancy.
Table 12: Partition robustness (LLaMA2-7B, 4-domain slice).
Variant Q&A (F1) NSum (ROUGE-L)
$\lambda=0$ (no synergy) 70.3 40.8
$\lambda=1.0$ (synergy-only) 70.2 40.6
[…]
Heuristic + Gaussian noise ($\sigma=0.05$) 70.2 40.7
Default (Table 11) 70.5 40.9
B.5 Scalability to many domains
We synthetically vary the number of domains $k$ and measure CPU partition overheads and peak GPU memory with 4-bit LoRA. Table 13 indicates sub-second CPU time and practical GPU usage up to k=50k{=}50. Partitioning is CPU-side and negligible relative to PEFT training; memory remains dominated by standard SFT/PEFT, confirming practicality at double-digit $k$.
Table 13: Large-$k$ partition costs and peak GPU memory (synthetic up to k=50k{=}50; A100).
$k$ Affinity build (time / RAM) Clustering time Peak GPU (4-bit LoRA)
20 0.47 s / 180 MB 0.14 s 19 GB
35 2.10 s / 620 MB 0.52 s 23 GB
50 3.20 s / 950 MB 0.90 s 26 GB
B.6 Scaling to twelve domains (real mixture)
We combine Wiki-10 (topic classification) and Multi-News (summarization), deduplicated to 12 domains, and keep the same hyper-parameters as the main study. Table 14 shows gains over all-in-one SFT and over a random 2-stage split while keeping memory low. The synergy-aware split generalizes beyond four domains to a heterogeneous, double-digit regime with consistent improvements.
Table 14: Twelve-domain mixture (LLaMA2-7B + LoRA).
Split strategy Avg. ACC (Wiki-10) ROUGE-L (Multi-News) Peak GPU (GB) Partition time (s)
All-in-one SFT 83.1 37.2 27.3 n/a
Random 2-stage 83.7 37.5 18.6 0.6
PMS-FTP (ours) 84.0 38.1 18.7 0.7
B.7 Inference footprint after LoRA merging
After each stage we merge the finished LoRA into the frozen backbone, so only one 4-bit adapter is carried at inference. Table 15 shows equal-or-lower memory than a single-adapter baseline. Together with the accuracy gains in Table 1, merging achieves a strictly better accuracy–memory trade-off than training/keeping multiple adapters.
Table 15: Measured inference memory (LLaMA2-7B, k=4k{=}4, A100).
Precision Tag-LLM (1 LoRA) PMS-FTP (merged) $\boldsymbol{\Delta}$
FP16 29.4 GB 28.7 GB $2.4\%$
INT8 19.1 GB 18.6 GB $2.6\%$
4-bit 17.2 GB 16.8 GB 0.4-0.4 GB
B.8 Empirical validity of the $\mathcal{G}$ objective
We sample 20 random partitions, compute $\mathcal{G}$, and measure worst-domain dev loss. Table 16 reports Pearson $\rho=-0.81$ ($p<0.01$), i.e., higher $\mathcal{G}$ predicts lower worst-domain error. This supports the practical usefulness of our bound-driven objective: $\mathcal{G}$ values correlate strongly with the metric we aim to improve.
Table 16: Correlation between $\mathcal{G}$ and worst-domain dev loss (LLaMA2-7B, 20 random partitions).
Statistic $\boldsymbol{\mathcal{G}}$ Worst-Dev Loss
Mean 0.432 1.72
Std 0.057 0.19
Min 0.318 1.38
Max 0.522 2.11
Pearson $\boldsymbol{\rho=-0.81}$ $(p<0.01,\;R^{2}\!\approx\!0.65)$
B.9 Reasoning benchmarks and reweighting baselines
[…]
Table 17: Reasoning tasks with 7B backbone (identical PEFT budgets).
Method Hellaswag (Acc) MMLU-STEM (Acc) ARC-easy (Acc) SciQ (Acc) GSM8K (Pass@1) Avg.
FULL 73.4 39.7 78.5 92.6 18.1 60.5
LoRA 73.1 39.3 77.9 92.4 17.5 60.0
Tag-LLM 74.8 41.1 79.6 93.3 19.3 61.6
iMTL 74.6 40.6 79.3 93.0 18.6 61.2
FAMO 73.7 41.3 78.6 92.6 18.7 60.8
ExcessMTL 74.2 40.5 79.7 92.1 17.5 60.8
PMS-FTP (ours) 75.6 42.1 80.6 93.8 19.9 62.4
Table 18: Reasoning tasks with 13B backbone (identical PEFT budgets).
Method Hellaswag (Acc) MMLU-STEM (Acc) ARC-easy (Acc) SciQ (Acc) GSM8K (Pass@1) Avg.
FULL 77.5 46.3 83.7 94.8 24.5 65.4
LoRA 77.1 45.9 83.2 94.6 23.8 64.9
Tag-LLM 79.0 47.3 84.7 95.4 25.1 66.3
iMTL 77.4 47.2 83.3 95.1 24.8 65.6
FAMO 77.7 47.6 84.5 95.2 24.2 65.8
ExcessMTL 78.1 47.0 83.1 95.0 23.8 65.4
PMS-FTP (ours) 79.9 49.0 85.7 95.8 26.8 67.4
B.10 Incremental addition of a new domain
After training on the original four domains, we add “Legal-QA” as a new stage and freeze prior adapters. Table 19 shows negligible forgetting on old domains and a gain over single-domain LoRA on the new domain. Disjoint, frozen adapters make PMS-FTP naturally amenable to one-shot domain extension without replay.
Table 19: One-shot incremental stage (add “Legal-QA”).
Model Avg. score on original 4 domains Legal-QA (EM)
Before add-on 65.5
After add-on (PMS-FTP) 65.4 68.2
Single-domain LoRA n/a 67.6
B.11 Stability of domain centroids
We (i) bootstrap mean SBERT embeddings to gauge centroid noise, and (ii) replace each single centroid by a K=3K{=}3 weighted barycenter. Table 20 shows bootstrap deviations are $3\%$ of the smallest inter-domain distance; Table 21 shows K=3K{=}3 centroids change final metrics by $\leq 0.1$ pp. A single mean embedding is a sufficiently stable domain signature for $\mathcal{G}$-guided clustering.
Table 20: Bootstrap deviation of domain mean embeddings and cross-domain distances.
Domain NN 95% CI of $\|\Delta\mu\|_{2}$ Cross-domain min dist
NSum 12,000 [0.038, 0.065] 1.92
Q&A 15,000 [0.031, 0.060] 2.04
Sent 10,500 [0.042, 0.072] 1.87
Topic 11,200 [0.040, 0.068] 1.99
Table 21: Single centroid vs. K=3K{=}3 weighted barycenters per domain.
Representation ROUGE-L (NSum) EM (Q&A) Sent (ACC) Topic (ACC) $\boldsymbol{\Delta}$ vs. 1-centroid
1 centroid (paper) 43.4 67.2 90.2 88.0
3 centroids (K=3) 43.3 67.1 90.1 87.9 0.1-0.1 $p$
Appendix C Proofs
C.1 Complete Proof for Theorem 3.1
Before formally beginning the proof, we first revisit the setting and present a key lemma:
There are $k$ source domains $\{\mathcal{D}_{j}\}_{j=1}^{k}$, each domain $\mathcal{D}_{j}$ with $n_{j}$ samples, total $n=\sum_{j=1}^{k}n_{j}$. We let $\alpha_{j}=\tfrac{n_{j}}{n}$ or any other nonnegative weighting such that $\sum_{j=1}^{k}\alpha_{j}=1$. Our LLM-based predictor $f_{\theta,\{\phi_{j}\}}$ is constrained so that
$\|\theta-\theta^{*}\|_{2}\;\leq\;\rho_{\theta},\quad\|\phi_{j}\|_{2}\;\leq\;\rho_{\phi},$ (18)
for each $j$. By Assumption 3.1, the model output is $L$-Lipschitz w.r.t. predictions, and the difference in outputs for different parameters is bounded by $B(\|\theta-\theta^{\prime}\|+\sum_{j}\|\phi_{j}-\phi_{j}^{\prime}\|)$. Thus the entire class of such $(\theta,\{\phi_{j}\})$ belongs to a low-capacity function family $\mathcal{F}$.
Lemma C.1.
Let $\mathcal{F}$ be the class of predictors with backbone/adapters bounded as above. Then for any $\delta\in(0,1)$, with probability at least $1-\delta$ over all $n$ samples,
$\bigl|\mathcal{L}^{\boldsymbol{\alpha}}(f)-\widehat{\mathcal{L}}^{\boldsymbol{\alpha}}(f)\bigr|\;\;\leq\;\;2LB\bigl(\rho_{\theta}+\sum_{j=1}^{k}\alpha_{j}\rho_{\phi}\bigr)\;+\;\sqrt{\tfrac{\ln(2/\delta)}{2n}}\qquad\forall\,f\in\mathcal{F},$ (19)
where $\widehat{\mathcal{L}}^{\boldsymbol{\alpha}}(f)=\sum_{j}\alpha_{j}\widehat{\mathcal{L}}_{\mathcal{D}_{j}}(f)$.
Proof.
Because the loss is $L$-Lipschitz and the model satisfies the output-difference bound from Assumption 3.1, each $f\in\mathcal{F}$ is $LB(\rho_{\theta}+\sum_{j}\rho_{\phi})$-Lipschitz in its parameters relative to $\ell_{2}$. Let $\mathcal{R}_{n}(\mathcal{F})$ be the empirical Rademacher complexity over the pooled sample of size $n$. Standard contraction (e.g. Mohri et al., 2018) yields
$\mathcal{R}_{n}(\mathcal{F})\;\leq\;LB\,(\rho_{\theta}+k\rho_{\phi})\,\frac{1}{\sqrt{n}}.$ (20)
Replacing $k\rho_{\phi}$ by $\sum_{j}\alpha_{j}\rho_{\phi}$ (because losses are weighted by $\alpha_{j}$) strengthens the constant. Applying the usual Rademacher tail bound with a union-bound over $\delta/2$ produces the stated inequality. ∎
Proof.
Here is the proof of Theorem 3.1.
Uniform convergence (empirical $\rightarrow$ true risk)
Define $\Gamma\bigl(\rho_{\theta},\rho_{\phi},\boldsymbol{\alpha},k\bigr):=2LB\bigl(\rho_{\theta}+\sum_{j}\alpha_{j}\rho_{\phi}\bigr).$ Lemma C.1 gives
$\mathcal{L}^{\boldsymbol{\alpha}}(f)\;\leq\;\widehat{\mathcal{L}}^{\boldsymbol{\alpha}}(f)\;+\;\Gamma(\rho_{\theta},\rho_{\phi},\boldsymbol{\alpha},k)\;+\;\sqrt{\tfrac{\ln(2/\delta)}{2n}}.$ (21)
Domain-discrepancy correction
Blitzer et al. [Blitzer, 2008] show that for any hypothesis $h$ and any two distributions $\mathcal{P},\mathcal{Q}$, $|\mathcal{L}_{\mathcal{P}}(h)-\mathcal{L}_{\mathcal{Q}}(h)|\leq d_{\mathcal{H}\!\Delta\!\mathcal{H}}(\mathcal{P},\mathcal{Q}).$ Summing over all pairs and using triangle inequality,
$\mathcal{L}^{\boldsymbol{\alpha}}(f)\;\leq\;\sum_{j=1}^{k}\alpha_{j}\mathcal{L}_{\mathcal{D}_{j}^{\!\!*}}(f)\;+\;\frac{1}{k}\sum_{i,j=1}^{k}d(\mathcal{D}_{i},\mathcal{D}_{j}),$ (22)
where $\mathcal{D}_{j}^{\!\!*}$ denotes drawing as if every example came from a single mixture domain-hence its empirical risk is exactly $\widehat{\mathcal{L}}^{\boldsymbol{\alpha}}(f)$, and the additive discrepancy penalty is weighted by $\beta:=1$ (absorbing the Lipschitz loss factor into the definition of $d$). Restoring the constant gives the $\beta$ appearing in the theorem.
Combine bounds
Insert (21) into (22):
$\mathcal{L}^{\boldsymbol{\alpha}}(f)\;\leq\;\widehat{\mathcal{L}}^{\boldsymbol{\alpha}}(f)\;+\;\Gamma(\rho_{\theta},\rho_{\phi},\boldsymbol{\alpha},k)\;+\;\frac{\beta}{k}\sum_{i,j=1}^{k}d(\mathcal{D}_{i},\mathcal{D}_{j})\;+\;\sqrt{\tfrac{\ln(2/\delta)}{2n}}.$
Replacing $\sqrt{\ln(2/\delta)/(2n)}$ by the big-$O(\sqrt{\ln(1/\delta)/n})$ notation and recalling $\widehat{\mathcal{L}}^{\boldsymbol{\alpha}}(f)=\sum_{j}\alpha_{j}\widehat{\mathcal{L}}_{\mathcal{D}_{j}}(f)$ yields exactly inequality (10), proving Theorem 3.1. ∎
C.2 Complete Proof for Theorem 3.2
Here is the proof of Theorem 3.2.
Proof.
For any fixed stage $t$ training on domains $S_{t}$, Theorem 3.1 with weights $\alpha_{j}^{t}:=\tfrac{n_{j}}{\sum_{i\in S_{t}}n_{i}}$ gives
$\sum_{j\in S_{t}}\alpha_{j}^{t}\,\mathcal{L}_{\mathcal{D}_{j}}(f^{t})$ $\leq\underbrace{\sum_{j\in S_{t}}\alpha_{j}^{t}\,\widehat{\mathcal{L}}_{\mathcal{D}_{j}}(f^{t})}_{\leq 1}+\;2LB\!\bigl(\rho_{\theta}+\rho_{\phi}\bigr)+\;\beta\,\!\!\!\sum_{\begin{subarray}{c}i,j\in S_{t}\\ i<j\end{subarray}}d(\mathcal{D}_{i},\mathcal{D}_{j})+\;O\!\bigl(\sqrt{\tfrac{\ln(1/\delta)}{\sum_{j\in S_{t}}n_{j}}}\bigr).$ (23)
Then we inject synergy and explicit capacity weight. Define  $\mathrm{Cap}(S_{t}):=\mu_{\theta}\|\Delta\theta^{t}\|_{2}^{2}+\mu_{\phi}\sum_{j\in S_{t}}\|\phi^{t}_{j}\|_{2}^{2}$. Because $\|\Delta\theta^{t}\|\!\leq\!\rho_{\theta}$ and $\|\phi_{j}^{t}\|\!\leq\!\rho_{\phi}$, we upper bound $2LB(\rho_{\theta}+\rho_{\phi})$ by $\mathrm{Cap}(S_{t})$ after tuning $\mu_{\theta},\mu_{\phi}$. Subtract and add $\lambda\!\sum_{i<j}s(\mathcal{D}_{i},\mathcal{D}_{j})$ to (23) to obtain
$\sum_{j\in S_{t}}\alpha_{j}^{t}\,\mathcal{L}_{\mathcal{D}_{j}}(f^{t})\;\leq\;1-\Bigl[-\!\!\!\sum_{i<j\in S_{t}}\!\!\bigl(d-\lambda s\bigr)-\mathrm{Cap}(S_{t})\Bigr]\;+\;O\!\bigl(\sqrt{\tfrac{\ln(1/\delta)}{N}}\bigr).$ (24)
Let $\mathcal{R}_{\max}(S_{1},\ldots,S_{M}):=\max_{t}{\textstyle\sum_{j\in S_{t}}\alpha_{j}^{t}\,\mathcal{L}_{\mathcal{D}_{j}}(f^{t})}.$ Taking the maximum of (24) over $t$ gives
$\mathcal{R}_{\max}(S_{1},\ldots,S_{M})\;\leq\;1-\mathcal{G}(S_{1},\ldots,S_{M})\;+\;O\!\bigl(\sqrt{\tfrac{\ln(1/\delta)}{N}}\bigr),$ (25)
because the bracketed term is exactly the $t$-th summand of $\mathcal{G}$ in (12). Define $\mathcal{B}(u):=[1-u]_{+}.$ Then $\mathcal{R}_{\max}\!\bigl(S_{1},\ldots,S_{M}\bigr)\leq\mathcal{B}(\mathcal{G}(S_{1},\ldots,S_{M}))+O(\sqrt{{\ln(1/\delta)}/{N}}).$
Because $\mathcal{B}$ is strictly decreasing on $(-\infty,1]$, maximising $\mathcal{G}$ minimises the bound. Hence the partition $(S_{1}^{*},\ldots,S_{M}^{*})$—the maximiser of $\mathcal{G}$—realises the smallest upper-bound, yielding (13). Any other split attains a weaker bound, completing the proof. ∎

### v3

@@
 1 Introduction
-Large language models (LLMs) have propelled natural language processing (NLP) to unprecedented capabilities (Kumar, 2024; Karanikolas et al., 2023; Hu et al., 2025; Zhang et al., 2025), owing largely to their extensive pretraining on massive, diverse textual corpora (Wu et al., 2022; Li et al., 2024a; Chen et al., 2025).
-Fine-tuning these pretrained models for a specific downstream domain has been widely explored and proven highly effective; notable approaches include adapter-based modules (Houlsby et al., 2019; Zhang et al., 2024a), parameter-efficient fine-tuning via low-rank updates (Hu et al., 2021), and instruction-based fine-tuning methods (Cao et al., 2024).
-These methods have also accelerated the application of LLMs in many scenarios, such as healthcare(Tong et al., 2025; Jiaqi Liu, 2025), transportation(Yao et al., 2023; Wang et al., 2025), and robotics(Xiao et al., 2025a; Li et al., 2025).
-However, while these methods excel in adapting to a single domain, practical scenarios frequently require simultaneous adaptation to multiple distinct domains-a scenario far less studied and substantially more challenging.
-Consider, for instance, a scenario where a single pretrained model must be simultaneously adapted to clinical texts (Thirunavukarasu et al., 2023), social media posts (Yang et al., 2024a; Jiang et al., 2024), and legal documents (Seabra et al., 2024).
-Naive approaches such as jointly fine-tuning the model across all domains or independently fine-tuning separate models per domain often yield suboptimal results(Xiao et al., 2025b): domain-specific features may negatively interfere, causing one domain to overshadow others or impair overall generalization capabilities (Lu et al., 2024; Zheng et al., 2024; Van Veen et al., 2023; Tan et al., 2025).
+Large language models (LLMs) have propelled natural language processing (NLP) to unprecedented capabilities (Kumar, 2024; Karanikolas et al., 2023; Hu et al., 2025; Zhang et al., 2025b; Lin et al., 2025), owing largely to their extensive pretraining on massive, diverse textual corpora (Wu et al., 2022; Li et al., 2024a; Chen et al., 2025b; Zhang et al., 2024a; Sun et al., 2025b).
+Fine-tuning these pretrained models for a specific downstream domain has been widely explored and proven highly effective; notable approaches include adapter-based modules (Houlsby et al., 2019; Zhang et al., 2024b), parameter-efficient fine-tuning via low-rank updates (Hu et al., 2021), and instruction-based fine-tuning methods (Cao et al., 2024).
+These methods have also accelerated the application of LLMs in many scenarios, such as healthcare(Tong et al., 2025; Liu et al., 2025; Wang et al., 2025), transportation(Yao et al., 2023; Lu et al., 2025b; Zeng et al., 2025b, a), robotics(Xiao et al., 2025a; Yan et al., 2025; Zhang et al., 2025a), and other applications (Sun et al., 2025a; Zhang et al., 2025c).
+However, while these methods excel in adapting to a single domain, practical scenarios frequently require simultaneous adaptation to multiple distinct domains-a scenario far less studied and substantially more challenging (Lu et al., 2025a; Chen et al., 2025a).
+Due to constraints on energy or computing power (He, 2025), we must consider how to enhance the model’s ability to adapt to different domains.
+Consider, for instance, a scenario where a single pretrained model must be simultaneously adapted to clinical texts (Thirunavukarasu et al., 2023), social media posts (Yang et al., 2024a; Jiang et al., 2024, 2025), and legal documents (Seabra et al., 2024).
+Naive approaches such as jointly fine-tuning the model across all domains or independently fine-tuning separate models per domain often yield suboptimal results(Xiao et al., 2025b; Zhang et al., 2023; Wang and Zhang, 2024; Wen et al., 2023): domain-specific features may negatively interfere, causing one domain to overshadow others or impair overall generalization capabilities (Lu et al., 2024a; Zheng et al., 2024; Van Veen et al., 2023; Tan et al., 2025).
 This motivates our central research question: how can we effectively and efficiently fine-tune a single LLM across multiple heterogeneous domains, exploiting inter-domain synergies while mitigating negative interference?
-A natural strategy to mitigate these challenges is to add adapter modules specialized for each domain (Houlsby et al., 2019; Zhang et al., 2024a).
+A natural strategy to mitigate these challenges is to add adapter modules specialized for each domain (Houlsby et al., 2019; Zhang et al., 2024b; Tao et al., 2023).
 While adapters reduce the need for full fine-tuning, they do not fully address the complexity of managing multiple source domains with potentially large discrepancies.
-In such cases, the learned model has to juggle both shared features (common linguistic properties across domains) and domain-specific features (rare words, styles, or content).
+In such cases, the learned model has to juggle both shared features (common linguistic properties across domains) and domain-specific features (rare words, styles, or content) (Lu et al., 2022).
 Existing approaches often handle these competing demands by imposing either domain adversarial objectives (Ganin and Lempitsky, 2015), distribution alignment (Peng et al., 2019), or low-rank parameter updates (Hu et al., 2021).
@@
 The success of LLMs such as GPT-3 (Brown et al., 2020), LLaMA (Touvron et al., 2023), and Falcon (Almazrouei et al., 2023) has underscored the importance of efficient fine-tuning.
-Fully updating model parameters incurs high computational costs (Zhang et al., 2024a; Hu et al., 2021), prompting parameter-efficient methods that modify only a small subset of parameters.
-Examples include adapter modules (Houlsby et al., 2019), low-rank projections (LoRA) (Hu et al., 2021), and tightly integrated adapters for LLaMA (Zhang et al., 2024a).
+Fully updating model parameters incurs high computational costs (Zhang et al., 2024b; Hu et al., 2021), prompting parameter-efficient methods that modify only a small subset of parameters.
+Examples include adapter modules (Houlsby et al., 2019), low-rank projections (LoRA) (Hu et al., 2021), and tightly integrated adapters for LLaMA (Zhang et al., 2024b).
 However, most methods assume a single domain; adapting LLMs efficiently to multiple domains remains challenging.
+Another related field is continual learning (Xu et al., 2025), but it assumes a sequential arrival of tasks, which differs from the setting in this paper.
 2.2 Multi-Domain Data Learning
-Real-world data often originate from disparate sources with distinct distributions and vocabularies (Zhang et al., 2024b; Ganin and Lempitsky, 2015; Sener and Koltun, 2018; Xiao, 2024; Xiao and Liu, 2025).
-Traditional multi-domain methods align representations through adversarial training (Pei et al., 2018; Ganin and Lempitsky, 2015), moment matching (Peng et al., 2019), or multi-task objectives (Royer et al., 2024), yet typically neglect inter-domain synergies.
+Real-world data often originate from disparate sources with distinct distributions and vocabularies (Ganin and Lempitsky, 2015; Sener and Koltun, 2018; Li et al., 2023; Zhang et al., 2024c; Chen et al., 2024; Xiao and Liu, 2025; Ke et al., 2025).
+Traditional multi-domain methods align representations through adversarial training (Pei et al., 2018; Ganin and Lempitsky, 2015), moment matching (Peng et al., 2019), or multi-task objectives (Royer et al., 2024; Shen and Zhang, 2025), yet typically neglect inter-domain synergies.
 Adapting LLMs further complicates this via memory overhead, forgetting pretrained knowledge, and cross-domain contamination.
@@
 Recent advances have introduced diverse metrics for data selection: instance-level criteria like perplexity (Cao et al., 2024), reward scores (Gou and Nguyen, 2024), and loss disparities (Li et al., 2024b), trajectory-based clustering via small proxy models (Yang et al., 2024b), as well as token-level selection methods (Lin et al., 2024).
-Existing works largely overlook one critical factor: domain interactions, where diversity metrics operate at instance/token levels (Pang et al., 2024) without modeling cross-domain compatibility.
+Existing works largely overlook one critical factor: domain interactions, where diversity metrics operate at instance/token levels (Pang et al., 2024) without modeling cross-domain compatibility (Li et al., 2025b).
 Our approach addresses this gap by systematically clustering domains for joint tuning.
@@
 Adapter (Houlsby et al., 2019) 41.3 64.1 88.9 86.3 42.3 65.7 89.5 87.2 42.9 67.0 90.2 88.0
-LLaMA-Adapter (Zhang et al., 2024a) 41.5 64.3 89.1 86.7 42.6 65.9 89.7 87.5 43.1 67.3 90.3 88.1
+LLaMA-Adapter (Zhang et al., 2024b) 41.5 64.3 89.1 86.7 42.6 65.9 89.7 87.5 43.1 67.3 90.3 88.1
 Q-LoRA (Dettmers et al., 2023) 41.7 64.4 89.0 86.5 42.4 65.6 89.5 87.3 43.0 67.2 90.1 87.9
@@
 Baselines
-We compare PMS-FTP against the following baselines: 1) Base Methods: Full Fine-Tuning (FULL), Fixed Backbone (FIXED); 2) Domain Adaptation: Multi-Domain Adversarial Network (MDAN) (Pei et al., 2018), Moment Matching (M3SDA) (Peng et al., 2019), Bayesian Gaussian Mixture (GMDI) (Ling et al., 2024); 3) Single-Domain LLM Fine-tuning: LoRA (Hu et al., 2021), Adapter (Houlsby et al., 2019), LLaMA-Adapter (Zhang et al., 2024a), Q-LoRA (Dettmers et al., 2023), Tag-LLM (Shen et al., 2024); 4) Data Selection: INSTRUCTMINING (IT) (Cao et al., 2024), S2L (Yang et al., 2024b).
+We compare PMS-FTP against the following baselines: 1) Base Methods: Full Fine-Tuning (FULL), Fixed Backbone (FIXED); 2) Domain Adaptation: Multi-Domain Adversarial Network (MDAN) (Pei et al., 2018), Moment Matching (M3SDA) (Peng et al., 2019), Bayesian Gaussian Mixture (GMDI) (Ling et al., 2024); 3) Single-Domain LLM Fine-tuning: LoRA (Hu et al., 2021), Adapter (Houlsby et al., 2019), LLaMA-Adapter (Zhang et al., 2024b), Q-LoRA (Dettmers et al., 2023), Tag-LLM (Shen et al., 2024); 4) Data Selection: INSTRUCTMINING (IT) (Cao et al., 2024), S2L (Yang et al., 2024b).
 Please refer to A.2 for more details on the baselines.
@@
 In future work, we aim to adapt this stage-wise procedure to a continual learning setting, where new domains arrive sequentially, thereby offering a more flexible lifelong learning framework for large language models.
+Furthermore, combining the proposed method with pruning strategies (Lu et al., 2024b; Zhou et al., 2025; Li et al., 2025a) represents an interesting direction, especially when dealing with models that have a massive number of parameters.
 Acknowledgements
@@
     • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The authors are encouraged to create a separate “Limitations” section in their paper.
     • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
@@
     • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The authors should answer “Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
     • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
@@
     • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
     • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
@@
     • Adapter [Houlsby et al., 2019]: Houlsby bottleneck adapters.
-    • LLaMA-Adapter [Zhang et al., 2024a]: zero-init residual adapters.
+    • LLaMA-Adapter [Zhang et al., 2024b]: zero-init residual adapters.
     • Q-LoRA [Dettmers et al., 2023]: 4-bit quantised LoRA layers.
@@
 $\mathrm{Syn}(\mathcal{D}_{i},\mathcal{D}_{j})=\tfrac{1}{2}\,\Bigl(\operatorname{Jacc}\bigl(V_{i},V_{j}\bigr)+\cos\bigl(\mu_{i},\mu_{j}\bigr)\Bigr),$ (17)
-where $V_{i}$ is the vocabulary set of $\mathcal{D}_{i}$, $\operatorname{Jacc}(V_{i},V_{j})=|V_{i}\cap V_{j}|/|V_{i}\cup V_{j}|$, and $\mu_{i}$ is the mean Sentence-BERT embedding of domain ii.
+where $V_{i}$ is the vocabulary set of $\mathcal{D}_{i}$, $\operatorname{Jacc}(V_{i},V_{j})=|V_{i}\cap V_{j}|/|V_{i}\cup V_{j}|$, and $\mu_{i}$ is the mean Sentence-BERT embedding of domain $i$.
 Both terms are in $[0,1]$; the average is therefore in $[0,1]$.
@@
 Table 15 shows equal-or-lower memory than a single-adapter baseline.
-Together with the accuracy gains in Table 1, merging achieves a strictly better accuracy–memory trade-off than training/keeping multiple adapters.
+Together with the accuracy gains in Table 1, merging achieves a strictly better accuracy-memory trade-off than training/keeping multiple adapters.
 Table 15: Measured inference memory (LLaMA2-7B, k=4k{=}4, A100).
@@
 B.10 Incremental addition of a new domain
-After training on the original four domains, we add “Legal-QA” as a new stage and freeze prior adapters.
+After training on the original four domains, we add Legal-QA as a new stage and freeze prior adapters.
 Table 19 shows negligible forgetting on old domains and a gain over single-domain LoRA on the new domain.
 Disjoint, frozen adapters make PMS-FTP naturally amenable to one-shot domain extension without replay.
-Table 19: One-shot incremental stage (add “Legal-QA”).
+Table 19: One-shot incremental stage (add Legal-QA).
 Model Avg. score on original 4 domains Legal-QA (EM)